# Rethinking Oversmoothing in Graph Neural Networks: A Rank-Based Perspective

### v2

@@ 6 Experiments
Empirical studies on the evolution of oversmoothing measures often use untrained, hundred-layer-deep GNNs (Rusch et al., 2022; Wang et al., 2022; Rusch et al., 2023b; Wu et al., 2023). We emphasize that this is an overly simplified setting. In more realistic settings, as the ones considered in this section, a trained GNN may suffer from significant performance degradation after only few-layers, at which stage the convergence patterns of most of the metrics are difficult to observe. The experiments that we present in this section validate the robustness of the effective rank and numerical rank in quantifying oversmoothing in GNNs against the other metrics.
In particular, we compare how different overmoothing metrics behave compared to the classification accuracy, varying the GNN architectures for node classification on real-world graph data. In our experiments, we consider the following metrics:
  • The Dirichlet Energy $E_{\mathrm{Dir}}$ (Cai & Wang, 2020; Rusch et al., 2023a) and its variant $E_{\mathrm{Proj}}$ (Wu et al., 2023). Both are discussed in Section 3.1, see in particular (8) and (9).
  • Normalized versions of the Dirichlet energy and its variant, namely $E_{\mathrm{Dir}}(X)/\|X\|_{F}^{2}$ and $E_{\mathrm{Proj}}(X)/\|X\|_{F}$. Indeed, from our previous discussion, a robust oversmoothing measure should be scale invariant with respect to the features. Metrics with global normalization like the ones we consider here have also been proposed in (Di Giovanni et al., 2023; Roth & Liebig, 2023; Maskey et al., 2023).
  • The Mean Average Distance (MAD) (Chen et al., 2020)
    $\text{MAD}(X)=\frac{1}{|\mathcal{E}|}\sum_{(i,j)\in\mathcal{E}}\left(1-\frac{X_{i}^{\top}X_{j}}{|X_{i}||X_{j}|}\right).$ (17)
@@
Figure 1: Four examples of GCNs corresponding to the first fours rows in table 2. For Erank and Numrank, we measure Erank$(X)-r^{*}_{\mathrm{ER}}$ and $\mathrm{NumRank}(X)-r^{*}_{\mathrm{NR}}$ for some $r^{*}>1$. In these particular cases, $r^{*}_{\mathrm{ER}}<1.85$, $r^{*}_{\mathrm{NR}}<1.3$. Note that the effective rank and numerical rank of the input features $X^{(0)}$ is about 1084 and 13.6, respectively.
In Table 2 and Figure 1, we train GNNs of a fixed hidden dimension equal to 32 on the Cora dataset. More results on Cora, as well as other datasets, are reported in Appendix B. We follow the standard setups of GCN and GAT as stated in equation (2) and (5), and use homogeneous LeakyReLU (LReLU) and subhomogenous Tanh as activation functions. For the GCN model, we also consider adding additional components, i.e. bias, LayerNorm (Ba et al., 2016) and PairNorm (Zhao & Akoglu, 2019). For each configuration, GNNs of eight different depths ranging from 2 to 24 are trained. The oversmoothing metric and accuracy results are averaged over 10 separately trained GNNs. All GNNs are trained with NAdam Optimizer and a constant learning rate of 0.01. The oversmoothing metrics are computed at the last hidden layer before the output layer. In Figure 1 and in Appendix B, we plot the behaviour of the different oversmoothing measures as well as the norm of the features and the accuracy of the trained GNNs when the depth of the network is increased. These figures clearly show that the network suffers a significant drop in accuracy, which is not matched by any visible change in standard oversmoothing metrics. By contrast, the rank of the feature representations decreases drastically, following quite closely the behaviour of the network’s accuracy. These findings are further sustained by the results shown in Table 2, where we compute the Pearson correlation coefficient between the logarithm of every measure and the classification accuracy of every GNN model. The use of a logarithmic transformation is based on the understanding that oversmoothing grows exponentially with the length of the network.
@@ 7 Conclusion
In this paper, we have discussed the problem of quantifying oversmoothing in GNNs. After discussing the limitations of the leading oversmoothing measures, we have proposed the use of the rank of the features as a better measure of oversmoothing. The experiments that we provided validate the robustness of the effective rank against the classical measures. In addition, we have proved theoretically the decay of the rank of the features for linear and nonnegative GNNs.
@@ Appendix A Proofs of the main results
A.1 proof of Lemma 3.2
Let $\pi:=\mathrm{span}\{uv^{T}\,|\;v\in\mathbb{R}^{d}\}$ be the $1$-dimensional matrix subspace of the rank-$1$ matrices having columns aligned to $u$. Then it is easy to note that given some matrix $X$, $(I-\mathcal{P})X$ provides the projection of the matrix $X$ on the subspace $\pi$, i.e.
@@
Indeed $\langle(I-\mathcal{P})X,uv^{T}\rangle_{F}=\mathrm{Tr}(vu^{T}(I-uu^{T})X)=0$. In particular, since the projection realizes the minimal distance, we have that
$\|X-\mathcal{P}X\|_{F}\leq\|X-uv^{T}\|_{F}\qquad\forall v\in\mathbb{R}^{d}.$ (19)
Now observe that $\sigma(uu^{T}A^{(l-1)}X^{(l-1)}W^{(l-1)})=u\bar{v}^{T}$ for some $\bar{v}$. Indeed, writing $v^{T}=u^{T}A^{(l-1)}X^{(l-1)}W^{(l-1)}$, we have that the $i$-th column of $\sigma(uu^{T}A^{(l-1)}X^{(l-1)}W^{(l-1)})$ is equal to $\sigma(v_{i}u)=\bar{v}_{i}u$ for some $\bar{v}_{i}$, because $u$ is an eigenvector of $\sigma$. As a consequence we have
$\|(I-\mathcal{P})X^{(l)}\|_{F}$ $\leq\|X^{(l)}-\sigma(uu^{T}A^{(l-1)}X^{(l-1)}W^{(l-1)})\|_{F}$ (20)
$=\|\sigma(A^{(l-1)}X^{(l-1)}W^{(l-1)})-\sigma(uu^{T}A^{(l-1)}X^{(l-1)}W^{(l-1)})\|_{F}$
@@
where we have used the $1$-Lipschitz property of $\sigma$.
@@
$\|\mathcal{P}X^{(l)}\|_{F}$ $\geq\|u^{T}A^{l}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}=$ (23)
$\|\big{(}\lambda_{1}^{l}v_{1}^{T}+u^{T}\widetilde{M}O\Big{(}{l\choose N}\lambda_{2}^{l-N}\Big{)}\widetilde{M^{\prime}}\big{)}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\geq$
$|\lambda_{1}|^{l}\Big{(}\|v_{1}^{T}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}-\Big{\|}u^{T}\widetilde{M}O\Big{(}{l\choose N}\Big{(}\frac{\lambda_{2}}{\lambda_{1}}\Big{)}^{l}\widetilde{M^{\prime}}X^{(0)}W^{(0)}\dots W^{(l-1)}\Big{\|}_{F}\Big{)}\geq$
$|\lambda_{1}|^{l}\|v_{1}^{T}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\Big{(}1-O\Big{(}{l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{T}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\Big{)}.$
Now observe that under the randomness hypothesis from (Furstenberg & Kifer, 1983) and more generally from the Oseledets ergodic multiplicative theorem, we have that for almost any $w\in\mathbb{R}^{d}$ the limit exists and is equal to the maximal Lyapunov exponent of the system, i.e. a constant $c(\nu)\geq 0$ depending only on the distribution $\nu$, $\lim_{l\rightarrow\infty}\frac{1}{l}\log\|w^{T}W^{(0)}\dots W^{(l-1)}\|=c(\nu)$. In particular for any $w$ and $\epsilon>0$ there exists $l_{w,\epsilon}$ sufficiently large such that for any $l>l_{w,\epsilon}$
$c(\nu)-\epsilon\leq\frac{1}{l}\log\|w^{T}W^{(0)}\dots W^{(l-1)}\|<c(\nu)+\epsilon.$ (24)
@@
$e^{l(c(\nu)-\epsilon)}\leq\|w^{T}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)}\quad\forall l\geq l_{w,\epsilon}$ (25)
Now take as vector $w$ first the rows of $X^{(0)}$ and then the vector $v_{1}^{T}X^{(0)}$, than almost surely for any $\epsilon$ there exists $l_{\epsilon}$ such that for any $l>l_{\epsilon}$
$e^{l(c(\nu)-\epsilon)}\leq\|w^{T}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)},$ (26)
holding for any $l\geq l_{\epsilon}$ and any $w\in\{v_{1}\}\cup\{X_{0}^{T}e_{i}\}_{i=1}^{N}$.
Next recall that $\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}=\sqrt{\sum_{i}\|e_{i}^{T}X^{(0)}W^{(0)}\|^{2}}$, meaning that almost surely, for $l\geq l_{\epsilon}$:
@@
($\Big{(}{l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{T}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\Big{)}\leq\Big{(}{l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}e^{2l\epsilon}\Big{)}$ (28)
@@
$\lim_{l}\frac{\|(I-\mathcal{P}X^{(l)})\|_{F}}{\|\mathcal{P}X^{(l)}\|_{F}}\sim\lim_{l}\frac{{l\choose N}|\lambda_{2}|^{l-N}\|X^{(0)}W^{(0)}\dots W^{(l)}\|_{F}}{|\lambda_{1}|^{l}\|v_{1}^{T}X^{(0)}W^{(0)}\dots W^{(l)}\|_{F}}=0$ (29)
@@ A.3 An illustrative example
By the contraction properties of the matrix $A$ we know that if $\max_{i}d_{H}\big{(}(AX)_{i},u\big{)}\leq C$, then
$d_{H}\Big{(}(AX)_{i},u\Big{)}=d_{H}\Big{(}(AX)_{i},\lambda_{1}(A)u\Big{)}=d_{H}\Big{(}(AX)_{i},Au\Big{)}\leq\beta d_{H}\big{(}X_{i},u\big{)}\qquad\forall i.$ (35)
@@
Then note that, for any $i$, we can write $F(X)_{i}$ as follows
$\big{(}F(X)\big{)}_{j}=\sum_{j}W_{ij}(AX)_{j}\,.$ (36)
@@
It remains to prove the claim. To this end, exploiting the expression of the Hilbert distance we write
@@
$\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{i}^{(l)},u)=0.$ (44)
Indeed by the uniform contractivity of $\{A^{(l)}\}$, if $C=\max_{i}d_{H}(X^{(0)}_{i},u)$, then there exists some $\beta_{C}<1$ such that
$\max_{i}d_{H}(X^{(l)}_{i},u)\leq\beta_{C}^{l}\max_{i}d_{H}(X^{(0)}_{i},u).$ (45)
@@
we have that (46) finally yields the thesis. To prove (46), we recall from Lemma 2.5.1 in (Lemmens & Nussbaum, 2012) that for any $w$ such that $u^{T}w=c$
@@
where $d_{T}(x,y)=\log(\max\{M(x/y),m(x/y)^{-1}\})$ and where since the dual cone of +$\mathbb{R}^{n}_{+}$ is +$\mathbb{R}^{n}_{+}$ itself, we are considering the norm induced by $u$ on the cone, i.e. $\|x\|_{u}=u^{T}x$ for any $x$ in the cone. In practice the norm induced by $u$ $\|\cdot\|_{u}$ can be defined by the Minkowki functional of the set $\Omega=\text{ConvexHull}\{\{\Omega_{1}\}\cup\{-\Omega_{1}\}\}$ where $\Omega_{1}=\{x\in\mathcal{K}\,\;u^{T}x\leq 1\}$
Then since $\|\mathcal{P}w\|_{u}=\|w\|_{u}=u^{T}w$, we have that $M(w/\mathcal{P}w)\geq 1$ and $m(w/\mathcal{P}w)\leq 1$. Thus $d_{T}(w,\mathcal{P}w)\leq d_{H}(w,\mathcal{P}w)$, yielding
@@
$\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}\leq C\|\mathcal{P}X^{(l)}\|_{F}^{2}\big{(}e^{\max_{i}\{d_{H}(X^{(l)}_{i},\mathcal{P}X^{(l)}_{i})\}}-1\big{)}^{2}.$ (51)
The proof is concluded observing that by the scale invariance property of the Hilbert distance $d_{H}(X^{(l)}_{i},\mathcal{P}X^{(l)}_{i})=d_{H}(X^{(l)}_{i},u)$, yielding:
$1\leq\big{(}\mathrm{NumRank}(X^{(l)})\big{)}\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|\mathcal{P}X^{(l)}\|_{F}^{2}}\leq 1+C\big{(}e^{\max_{i}\{d_{H}(X^{(l)}_{i},u)\}}-1\big{)}^{2}.$ (52)

### v3

@@ 6 Experiments
Empirical studies on the evolution of oversmoothing measures often use untrained, hundred-layer-deep GNNs [35, 44, 34, 45]. We emphasize that this is an overly simplified setting. In more realistic settings, as the ones considered in this section, a trained GNN may suffer from significant performance degradation after only few-layers, at which stage the convergence patterns of most of the metrics are difficult to observe. The experiments that we present in this section validate the robustness of the relaxed rank metrics in quantifying oversmoothing in GNNs against the other metrics.
In particular, we compare how different oversmoothing metrics behave compared to the classification accuracy, varying the GNN architectures for node classification on real-world graph data. In our experiments, we consider the following metrics:
  • The Dirichlet Energy $E_{\mathrm{Dir}}$ [4, 33] and its variant $E_{\mathrm{Proj}}$ [45]. Both are discussed in Section 3.1, see in particular (1) and (2).
  • Normalized versions of Dirichlet energy and its variant, $E_{\mathrm{Dir}}(X)/\|X\|_{F}^{2}$ and $E_{\mathrm{Proj}}(X)/\|X\|_{F}$. Indeed, from our previous discussion, a robust oversmoothing measure should be scale invariant with respect to the features. Metrics with global normalization like the ones we consider here have also been proposed in [8, 30, 20].
  • The Mean Average Distance (MAD) [5]
    $\textstyle\text{MAD}(X)=\frac{1}{|\mathcal{E}|}\sum_{(i,j)\in\mathcal{E}}\Big{(}1-\frac{X_{i,:}^{\top}X_{j,:}}{|X_{i,:}||X_{j,:}|}\Big{)}.$
@@
Figure 2: Four examples of the metric behaviours computed at the last hidden layer of separately trained GCNs of increasing depths. For Erank and Numrank, we measure Erank$(X)-r^{*}_{\mathrm{ER}}$ and $\mathrm{NumRank}(X)-r^{*}_{\mathrm{NR}}$ for some $r^{*}>1$. In these particular cases, $r^{*}_{\mathrm{ER}}<1.85$, $r^{*}_{\mathrm{NR}}<1.3$. Note that the effective rank and numerical rank of the input features $X^{(0)}$ are about 1084 and 13.6, respectively. Additional results are attached in Appendix G.
In Table 1 and Figure 2, we train GNNs of a fixed hidden dimension equal to 32 on homophilic, heterophilic and large-scale datasets in their default splits. We follow the standard setups of GCN and GAT as stated in Sections 2.1 and 2.2, and use homogeneous LeakyReLU (LReLU) as the activation function. For each configuration, GNNs of eight different depths ranging from 2 to 24 are trained. The oversmoothing metric and accuracy results are averaged over 10 separately trained GNNs. All GNNs are trained with NAdam Optimizer and a constant learning rate of 0.01. The oversmoothing metrics are computed at the last hidden layer before the output layer. In Figure 2 and in Appendix G, we plot the behaviour of the different oversmoothing measures, the norm of the features, and the accuracy of the trained GNNs with increasing depth. These figures clearly show that the network suffers a significant drop in accuracy, which is not matched by any visible change in standard oversmoothing metrics. By contrast, the rank of the feature representations decreases drastically, following closely the behaviour of the network’s accuracy. These findings are further sustained by the results shown in Table 1, where we compute the Pearson correlation coefficient between the logarithm of every measure and the classification accuracy of every GNN model. The use of a logarithmic transformation is based on the understanding that oversmoothing grows exponentially with the length of the network.
Extensions of these results are provided in Appendices E and F. In Appendix D, we perform an asymptotic ablation study on very deep (300-layer) synthetic networks with randomly sampled, untrained weights. This study serves to validate our theoretical findings on the convergence of relaxed rank metrics and to demonstrate that such untrained settings offer little insight into the ability of existing metrics to quantify oversmoothing in realistic, trained networks.
@@ 7 Conclusion
In this paper, we have discussed the problem of quantifying oversmoothing in message-passing GNNs. After simplifying the existing theoretical analysis using nonlinear activation eigenvectors and discussing the limitations of the leading oversmoothing measures, we propose the use of the rank of the features as a better measure of oversmoothing. We provide extensive experiments to validate the robustness of the effective rank against the classical measures. In addition, we have proved theoretically the decay of the rank of the features for message-passing GNNs.
Acknowledgement
KZ was supported by the EPSRC Centre for Doctoral Training in Mathematical Modelling, Analysis and Computation (MAC-MIGS) funded by the UK Engineering and Physical Sciences Research Council (grant EP/S023291/1), Heriot-Watt University and the University of Edinburgh. PD and FT are members of the Gruppo Nazionale Calcolo Scientifico - Istituto Nazionale di Alta Matematica (GNCS-INdAM). PD was supported by the MUR-PRO3 grant STANDS. DJH was supported by a Fellowship from the Leverhulme Trust. FT was partially funded by the PRIN-MUR project MOLE, code 2022ZK5ME7, and PRIN-PNRR project FIN4GEO within the European Union’s Next Generation EU framework, Mission 4, Component 2, CUP P2022BNB97.
@@ Appendix A Proofs of the main results
A.1 Proof of Theorem 3.2
We start proving that
$\|(I-\mathcal{P})X^{(l+1)}\|_{F}\leq\|(I-\mathcal{P})A^{(l)}X^{(l)}W^{(l)}\|_{F},$ (6)
where $\mathcal{P}=uu^{\top}/\|u\|^{2}$ is the projection matrix on the linear space spanned by $u$.
To this end, let $\pi:=\mathrm{span}\{uv^{\top}\,|\;v\in\mathbb{R}^{d}\}$ be the $1$-dimensional matrix subspace of the rank-$1$ matrices having columns aligned to $u$. Then it is easy to note that given some matrix $X$, $(I-\mathcal{P})X$ provides the projection of the matrix $X$ on the subspace $\pi$, i.e.
@@
Indeed $\langle(I-\mathcal{P})X,uv^{\top}\rangle_{F}=\mathrm{Tr}(vu^{\top}(I-uu^{\top}/\|u\|^{2})X)=0$. In particular, since the projection realizes the minimal distance, we have that
$\|X-\mathcal{P}X\|_{F}\leq\|X-uv^{\top}\|_{F}\qquad\forall v\in\mathbb{R}^{d}.$ (8)
Now observe that $\sigma(uu^{\top}A^{(l-1)}X^{(l-1)}W^{(l-1)})=u\bar{v}^{\top}$ for some $\bar{v}$. Indeed, writing $v^{\top}=u^{\top}A^{(l-1)}X^{(l-1)}W^{(l-1)}$, we have that the $i$-th column of $\sigma(uu^{\top}A^{(l-1)}X^{(l-1)}W^{(l-1)})$ is equal to $\sigma(v_{i}u)=\bar{v}_{i}u$ for some $\bar{v}_{i}$, because $u$ is an eigenvector of $\sigma$. As a consequence we have
$\|(I-\mathcal{P})X^{(l)}\|_{F}$ $\leq\|X^{(l)}-\sigma(uu^{\top}A^{(l-1)}X^{(l-1)}W^{(l-1)})\|_{F}$ (9)
$=\|\sigma(A^{(l-1)}X^{(l-1)}W^{(l-1)})-\sigma(uu^{\top}A^{(l-1)}X^{(l-1)}W^{(l-1)})\|_{F}$
@@
where we have used the $1$-Lipschitz property of $\sigma$. This concludes the proof of (6)
To conclude the proof of the theorem observe that, in the decomposition
$(I-\mathcal{P})A^{(l)}=(I-\mathcal{P})A^{(l)}\mathcal{P}+(I-\mathcal{P})A^{(l)}(I-\mathcal{P}),$
the matrix $(I-\mathcal{P})A^{(l)}\mathcal{P}$ is zero because $A^{(l)}u=\lambda_{1}^{l}u$ for any $l$. Thus
$(I-\mathcal{P})A^{(l)}=(I-\mathcal{P})A^{(l)}(I-\mathcal{P}),$
and, from (6) and the inequality $\|AB\|_{F}\leq\|A\|_{2}\|B\|_{F}$, we have
$\|(I-\mathcal{P})X^{(L)}\|_{F}\leq\Big{(}\Pi_{l=0}^{L-1}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}\Big{)}\|X^{(0)}\|_{F}.$
So the thesis follows from the hypothesis about the product $\Pi_{l=0}^{L-1}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}$.
@@
$\|\mathcal{P}X^{(l)}\|_{F}\geq\|u^{\top}A^{l}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}=$ (12)
$=\|\big{(}\lambda_{1}^{l}v_{1}^{\top}+u^{\top}\widetilde{M}O\Big{(}{l\choose N}\lambda_{2}^{l-N}\Big{)}\widetilde{M^{\prime}}\big{)}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\geq$
$\geq|\lambda_{1}|^{l}\Big{(}\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}-\Big{\|}u^{\top}\widetilde{M}O\Big{(}{l\choose N}\Big{(}\frac{\lambda_{2}}{\lambda_{1}}\Big{)}^{l}\widetilde{M^{\prime}}X^{(0)}W^{(0)}\dots W^{(l-1)}\Big{\|}_{F}\Big{)}\geq$
$\geq|\lambda_{1}|^{l}\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\Big{(}1-O\Big{(}{l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\Big{)}$
Now observe that under the randomness hypothesis from [11] and more generally from the Oseledets ergodic multiplicative theorem, we have that for almost any the limit $w\in\mathbb{R}^{d}$ $\lim_{l\rightarrow\infty}\frac{1}{l}\log\|w^{\top}W^{(0)}\dots W^{(l-1)}\|=c(\nu)$ exists and is equal to the maximal Lyapunov exponent of the system. In particular for any $w$ and $\epsilon>0$ there exists $l_{w,\epsilon}$ sufficiently large such that for any $l>l_{w,\epsilon}$
$c(\nu)-\epsilon\leq\frac{1}{l}\log\|w^{\top}W^{(0)}\dots W^{(l-1)}\|<c(\nu)+\epsilon$ (13)
@@
$e^{l(c(\nu)-\epsilon)}\leq\|w^{\top}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)}\quad\forall l\geq l_{w,\epsilon}$ (14)
Now take as vector $w$ first the rows of $X^{(0)}$ and then the vector $v_{1}^{\top}X^{(0)}$, than almost surely for any $\epsilon$ there exists $l_{\epsilon}$ such that for any $l>l_{\epsilon}$
$e^{l(c(\nu)-\epsilon)}\leq\|w^{\top}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)},$ (15)
holding for any $l\geq l_{\epsilon}$ and any $w\in\{v_{1}\}\cup\{X_{0}^{\top}e_{i}\}_{i=1}^{N}$.
Next recall that $\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}=\sqrt{\sum_{i}\|e_{i}^{\top}X^{(0)}W^{(0)}\|^{2}}$, meaning that almost surely, for $l\geq l_{\epsilon}$:
@@
($\left({l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\right)\leq\left({l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}e^{2l\epsilon}\right)$ (17)
@@
$\lim_{l}\frac{\|(I-\mathcal{P}X^{(l)})\|_{F}}{\|\mathcal{P}X^{(l)}\|_{F}}\sim\lim_{l}\frac{{l\choose N}|\lambda_{2}|^{l-N}\|X^{(0)}W^{(0)}\dots W^{(l)}\|_{F}}{|\lambda_{1}|^{l}\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l)}\|_{F}}=0$ (18)
@@ A.3 An illustrative example
By the contraction properties of the matrix $A$ we know that if $\max_{i}d_{H}\big{(}(AX)_{:,i},u\big{)}\leq C$, then
$d_{H}\Big{(}(AX)_{:,i},u\Big{)}=d_{H}\Big{(}(AX)_{:,i},\lambda_{1}(A)u\Big{)}=d_{H}\Big{(}(AX)_{:,i},Au\Big{)}\leq\beta d_{H}\big{(}X_{:,i},u\big{)}\qquad\forall i.$ (24)
@@
Then note that, for any $i$, we can write $F(X)_{:,j}$ as follows
$\big{(}F(X)\big{)}_{:,j}=\sum_{j}W_{ij}(AX)_{:,j}\,.$ (25)
@@
We only miss to prove the claim. To this end, exploiting the expression of the Hilbert distance we write
@@
$\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{:,i}^{(l)},u)=0.$ (33)
Indeed by the uniform contractivity of $\{A^{(l)}\}$, if $C=\max_{i}d_{H}(X^{(0)}_{:,i},u)$, then there exists some $\beta_{C}<1$ such that
$\max_{i}d_{H}(X^{(l)}_{:,i},u)\leq\beta_{C}^{l}\max_{i}d_{H}(X^{(0)}_{:,i},u).$ (34)
@@
we have that (35) finally yields the thesis. To prove (35), we recall from Lemma 2.5.1 in [18] that for any $w$ such that $u^{\top}w=c$
@@
where $d_{T}(x,y)=\log(\max\{M(x/y),m(x/y)^{-1}\})$ and where since the dual cone of +$\mathbb{R}^{n}_{+}$ is +$\mathbb{R}^{n}_{+}$ itself, we are considering the norm induced by $u$ on the cone, i.e. $\|x\|_{u}=u^{\top}x$ for any $x$ in the cone. In practice the norm induce by $u$ $\|\cdot\|_{u}$ can be defines by the Minkowki functional of the set $\Omega=\text{ConvexHull}\{\{\Omega_{1}\}\cup\{-\Omega_{1}\}\}$ where $\Omega_{1}=\{x\in\mathcal{K}\,\;u^{\top}x\leq 1\}$
Then since $\|\mathcal{P}w\|_{u}=\|w\|_{u}=u^{\top}w$, we have that $M(w/\mathcal{P}w)\geq 1$ and $m(w/\mathcal{P}w)\leq 1$. Thus $d_{T}(w,\mathcal{P}w)\leq d_{H}(w,\mathcal{P}w)$, yielding
@@
$\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}\leq C\|\mathcal{P}X^{(l)}\|_{F}^{2}\big{(}e^{\max_{i}\{d_{H}(X^{(l)}_{:,i},\mathcal{P}X^{(l)}_{:,i})\}}-1\big{)}^{2}.$ (40)
the proof is concluded using (33) and observing that by the scale invariance property of the Hilbert distance $d_{H}(X^{(l)}_{:,i},\mathcal{P}X^{(l)}_{:,i})=d_{H}(X^{(l)}_{:,i},u)$, yielding:
$1\leq\big{(}\mathrm{NumRank}(X^{(l)})\big{)}\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|\mathcal{P}X^{(l)}\|_{F}^{2}}\leq 1+C\big{(}e^{\max_{i}\{d_{H}(X^{(l)}_{:,i},u)\}}-1\big{)}^{2}$ (41)
and concluding the proof.